# on CLIP via Concept Confusion

### v1

@@ A.5 Implementation Details
In our main experiments, we use the CIFAR-10, CIFAR-100, and ImageNet-Tiny datasets. The image encoder is derived from CLIP ViT B/16, and we employ TCAV [22] as the concept extractor. Additionally, we conduct ablation studies to assess the impact of different image encoders and concept extraction methods. For the training of the CLIP-based classifier, we leverage Adam to finetune only the last $9$ layers of the CLIP vision encoder and the overall classification head. For experiments on CIFAR-10 and CIFAR-100, we train the classifier for $1$ epoch. For experiments on Tiny-ImageNet, we train the classifier for $3$ epochs. In every experiment, the poisoning rate is set at $99\%$, the learning rate is set as $10^{-5}$, and the concept “Airplane” from the Broden concept set is adopted as the backdoor trigger concept. Results are reported based on four repeated experiment runs.
Appendix B Ablation Study
B.1 Impact of Various Encoder Architectures
@@
We evaluated our attack methodology on the CIFAR-10 dataset across four distinct CLIP-ViT architectures, utilizing the ”Airplane” concept as the trigger and the corresponding ”Airplane” class as the target label. The results, presented in Tab. 8, demonstrate remarkable consistency with perfect Attack Success Rates (ASR) of 100% and high CACC maintained across all tested architectures. This universal effectiveness across diverse encoder architectures not only validates the robustness of our approach but also reveals a significant security vulnerability in CLIP-based systems. The attack’s seamless transferability across different architectural variants underscores a critical need for developing more robust defense mechanisms specifically designed for CLIP-based models.
B.2 Impact of Poisons Rates
@@
We investigated the relationship between poisoned data ratios and attack efficacy by conducting experiments on the CIFAR-10 dataset, designating ”Airplane” as the target label and employing three distinct concepts: ”Airplane,” ”Engine,” and ”Headlight.” The results, documented in Tab. 8, demonstrate remarkable attack resilience across varying poisoning ratios. Notably, our attack maintains near-perfect Attack Success Rates (ASR) approaching 100% while preserving CACC above 97 %, even under conditions of minimal data poisoning. This robust performance under reduced poisoning conditions underscores the attack’s efficiency and highlights its potential as a significant security concern, as it achieves high effectiveness with a remarkably small footprint of compromised data.
@@
B.3 Impact of Various Concepts
The concept ablation experiment is conducted under CIFAR-10 using TCAV [22] as the Concept Extractor on the CIFAR-10 dataset and CLIP-ViT-B/16. With our method, we apply backdoor attack on 30 different concepts. The results are shown in Tab. 9.
[…]
@@ B.3 Impact of Various Concepts
Appendix C Concept Extractor
C.1 TCAV
TCAV [22] is an important method for obtaining interpretable concepts in machine learning models. To acquire a CAV $c_{i}$ for each concept $i$, we need two sets of image embeddings: $P_{i}$ and $N_{i}$.
$P_{i}$ =$=\{f(x^{p}_{1}),\ldots,f(x^{p}_{N_{p}})\}$
$N_{i}$ =$=\{f(x^{n}_{1}),\ldots,f(x^{n}_{N_{n}})\}$
@@ C.1 TCAV
C.2 Label-free Concept Bottleneck Models
Label-free concept bottleneck models (Label-free CBM, Oikarinen et al. [40]) can transform any neural network into an interpretable concept bottleneck model without the need for concept-annotated data while maintaining the task accuracy of the original model, which significantly saves human and material resources.
@@ C.2 Label-free Concept Bottleneck Models
A. Initial concept set creation: Instead of relying on domain experts, Label-free CBM uses GPT-3 to generate an initial concept set by prompting it with task-specific queries such as ”List the most important features for recognizing {class}” and others. Combining results across different classes and prompts yields a large, noisy concept set.
@@
where $R_{\alpha}(W_{F})=(1-\alpha)\frac{1}{2}||W_{F}||_{F}+\alpha||W_{F}||_{1,1}$, and $\lambda$ controls the level of sparsity. The GLM-SAGA solver is used to optimize this step, and $\alpha=0.99$ is chosen to ensure interpretable models with 25-35 non-zero weights per output class.
C.3 Semi-supervised Concept Bottleneck Models
By leveraging joint training on both labeled and unlabeled data and aligning the unlabeled data at the conceptual level, semi-supervised concept bottleneck models (Semi-supervised CBM, Hu et al. [20]) address the challenge of acquiring large-scale concept-labeled data in real-world scenarios. Their approach can be summarized as follows:
@@ C.3 Semi-supervised Concept Bottleneck Models
$\mathcal{H}_{p,q,i}=\frac{\bm{e}_{i}^{\top}V_{p,q}}{||\bm{e}_{i}||\cdot||V_{p,q}||},\quad p=1,\ldots,H,\quad q=1,\ldots,W,$
@@
Appendix D Detection Experiment
We train 10 backdoored models, each using a different concept, and evaluate their detection accuracy under $C^{2}$ttack. Tab. 10 presents the overall detection accuracy, while Tab. 11 provides detailed detection results for each backdoored model. “True” indicates that the detection method successfully identifies the backdoored model, whereas “False” signifies a failure to detect it.
Table 10: Detection accuracy against $C^{2}$ttack. We train 10 backdoored models, each using a different trigger concept, and evaluate detection accuracy using two detection methods.
@@ Appendix D Detection Experiment
Table 11: Detailed detection results for each backdoored model. “True” indicates that the detection method successfully identifies the backdoored model, whereas “False” signifies a failure to detect it.

### v2

@@ B.5 Implementation Details
In our main experiments, we use the CIFAR-10, CIFAR-100, and ImageNet-Tiny datasets. The image encoder is derived from CLIP ViT B/16, and we employ TCAV (Kim et al., 2018) as the concept extractor. Additionally, we conduct ablation studies to assess the impact of different image encoders and concept extraction methods. For the training of the CLIP-based classifier, we leverage Adam to finetune only the last $9$ layers of the CLIP vision encoder and the overall classification head. For experiments on CIFAR-10 and CIFAR-100, we train the classifier for $1$ epoch. For experiments on Tiny-ImageNet, we train the classifier for $3$ epochs. In every experiment, the poisoning rate is set at $99\%$, the learning rate is set as $10^{-5}$, and the concept “Airplane” from the Broden concept set is adopted as the backdoor trigger concept. Results are reported based on four repeated experiment runs.
Appendix C Ablation Study
C.1 Impact of Various Encoder Architectures
@@
We evaluated our attack methodology on the CIFAR-10 dataset across four distinct CLIP-ViT architectures, utilizing the "Airplane" concept as the trigger and the corresponding "Airplane" class as the target label. The results, presented in Tab. 7, demonstrate remarkable consistency with perfect Attack Success Rates (ASR) of 100% and high CACC maintained across all tested architectures. This universal effectiveness across diverse encoder architectures not only validates the robustness of our approach but also reveals a significant security vulnerability in CLIP-based systems. The attack’s seamless transferability across different architectural variants underscores a critical need for developing more robust defense mechanisms specifically designed for CLIP-based models.
C.2 Impact of Poisons Rates
@@
We investigated the relationship between poisoned data ratios and attack efficacy by conducting experiments on the CIFAR-10 dataset, designating "Airplane" as the target label and employing three distinct concepts: "Airplane," "Engine," and "Headlight." The results, documented in Tab. 7, demonstrate remarkable attack resilience across varying poisoning ratios. Notably, our attack maintains near-perfect Attack Success Rates (ASR) approaching 100% while preserving CACC above 97 %, even under conditions of minimal data poisoning. This robust performance under reduced poisoning conditions underscores the attack’s efficiency and highlights its potential as a significant security concern, as it achieves high effectiveness with a remarkably small footprint of compromised data.
@@
C.3 Impact of Various Concepts
The concept ablation experiment is conducted under CIFAR-10 using TCAV (Kim et al., 2018) as the Concept Extractor on the CIFAR-10 dataset and CLIP-ViT-B/16. With our method, we apply backdoor attack on 30 different concepts. The results are shown in Tab. 8.
Table 8: Clean Accuracy (CACC) ($\%$) and Attack Sucess Rate (ASR) ($\%$) of different concepts.
@@ C.3 Impact of Various Concepts
TCAV Label-free Semi-supervise
Concept CACC ASR CACC ASR CACC ASR
Airplane 97.8 100 97.2 100 97.6 100
Oven 97.6 100 96.8 100 97.6 100
Engine 97.5 100 97.3 100 96.8 100
Headlight 97.2 100 97.3 100 97.2 97.7
Head 97.2 100 97.3 97.0 97.1 100
Clock 97.1 100 96.8 100 97.4 100
Mirror 97.0 100 96.7 100 95.9 100
Air-conditioner 97.0 100 97.4 100 97.4 100
Building 96.5 100 97.0 100 96.9 95.7
Cushion 96.4 100 97.4 95.7 97.2 98.6
Appendix D Concept Extractor
D.1 TCAV
TCAV (Kim et al., 2018) is an important method for obtaining interpretable concepts in machine learning models. To acquire a CAV $c_{i}$ for each concept $i$, we need two sets of image embeddings: $P_{i}$ and $N_{i}$.
$P_{i}$ $=\{f(x^{p}_{1}),\ldots,f(x^{p}_{N_{p}})\}$
$N_{i}$ $=\{f(x^{n}_{1}),\ldots,f(x^{n}_{N_{n}})\}$
@@ D.1 TCAV
D.2 Label-free Concept Bottleneck Models
Label-free concept bottleneck models (Label-free CBM (Oikarinen et al., 2023)) can transform any neural network into an interpretable concept bottleneck model without the need for concept-annotated data while maintaining the task accuracy of the original model, which significantly saves human and material resources.
@@ D.2 Label-free Concept Bottleneck Models
A. Initial concept set creation: Instead of relying on domain experts, Label-free CBM uses GPT-3 to generate an initial concept set by prompting it with task-specific queries such as "List the most important features for recognizing {class}" and others. Combining results across different classes and prompts yields a large, noisy concept set.
@@
[…]
D.3 Semi-supervised Concept Bottleneck Models
By leveraging joint training on both labeled and unlabeled data and aligning the unlabeled data at the conceptual level, semi-supervised concept bottleneck models (Semi-supervised CBM (Hu et al., 2024)) address the challenge of acquiring large-scale concept-labeled data in real-world scenarios. Their approach can be summarized as follows:
@@ D.3 Semi-supervised Concept Bottleneck Models
[…]
@@
Appendix E Detection Experiment
We train 10 backdoored models, each using a different concept, and evaluate their detection accuracy under $C^{2}$Attack. Tab. 10 presents the overall detection accuracy, while Tab. 11 provides detailed detection results for each backdoored model. “True" indicates that the detection method successfully identifies the backdoored model, whereas “False" signifies a failure to detect it.
Table 10: Detection accuracy against $C^{2}$Attack. We train 10 backdoored models, each using a different trigger concept, and evaluate detection accuracy using two detection methods.
@@ Appendix E Detection Experiment
Table 11: Detailed detection results for each backdoored model. “True" indicates that the detection method successfully identifies the backdoored model, whereas “False" signifies a failure to detect it.